1.   Assuming that an oscilloscope displays a Vmax of 5.22V and Vmin of 1.33V. Calculate the percentage of modulation. 



1
Expert's answer
2021-10-01T03:57:55-0400

Given "V_{max}=5.22V" and "V_{min}=1.33V," modulation(m) is given by,

"m=((V_{max}-V_{min})\/(V_{max}+V_{min}))"

"m=(5.22-1.33)\/(5.22+1.33)"

"m=3.89/6.55=0.5939(4\displaystyle\space decimal\displaystyle\space places)"

Therefore, percentage of modulation is given as,

"percentage\displaystyle\space of\displaystyle\space modulation=m*100=0.5939*100=59.39\%"